# GradTree: Learning Axis-Aligned Decision Trees with Gradient Descent

### v6

@@ 2 Related Work
While optimal DTs search for a global optimum, GradTree does not necessarily pursue this. Instead, like optimization in neural networks, it aims for a local optimum that offers good generalization to test data. We further want to emphasize that the local optima that can be reached by GradTree have a significant advantage over the local optimum of a greedy approach: While the local optimum of greedy approaches is constrained by sequentially selecting the optimal split at each node, GradTree overcomes this limitation by optimizing all parameters jointly.
@@ 3.2 Dense Decision Tree Representation
To this end, we expand the vector $\boldsymbol{\iota}\in\mathbb{R}^{2^{d}-1}$ to a matrix $I\in\mathbb{R}^{2^{d}-1}\times\mathbb{R}^{n}$. This is achieved by one-hot encoding the feature index as $\boldsymbol{\iota}\in\mathbb{R}^{n}$ for each internal node. This adjustment is necessary for the optimization process to account for the fact that feature indices are categorical instead of ordinal. Although our matrix representation for feature selection has parallels with that proposed by Popov, Morozov, and Babenko (2019), we introduce a novel aspect: A matrix representation for split thresholds. We denote this representation as $T\in\mathbb{R}^{2^{d}-1}\times\mathbb{R}^{n}$. Instead of representing a single value for all features, we store individual values for each feature, denoted as $\boldsymbol{\tau}\in\mathbb{R}^{n}$. This modification is tailored to support the optimization process, particularly in recognizing that split thresholds are feature-specific and non-interchangeable. In essence, a viable split threshold for one feature may not be suitable for another. This adjustment acts as a memory mechanism, ensuring that a given split threshold is exclusively associated with the corresponding feature. Consequently, this refinement enhances the exploration of feature selection at every split during the training.
Besides the previously mentioned advantages, using a dense DT representation allows the use of matrix multiplications for an efficient computation. Accordingly, we can reformulate the Heaviside split function (Equation 2) as
@@ 4.2 Results
Binary
Multi
Binary
Multi
Binary Multi
GradTree
054 086
0.051
0.174
0.670 0.470
DNDT
887 907
0.039
0.239
0.306 0.295
GeneticTree
007 020
0.204
0.258
0.427 0.315
DL8.5
028 029
0.202
0.260
0.371 0.331
CART
067 076
0.183
0.247
0.571 0.929
@@
ST Entmax (ours)
ST Gumbel
Temp. Annealing
@@ Appendix D Datasets
Blood Transfusion
4
748
178 / 570
2
https://archive.ics.uci.edu/ml/datasets/Blood+Transfusion+Service+Center
Banknote Authentication
4
1372
610 / 762
2
https://archive.ics.uci.edu/ml/datasets/banknote+authentication
Titanic
7
891
342 / 549
2
https://www.kaggle.com/c/titanic
Raisin
7
900
450 /450
2
https://archive.ics.uci.edu/ml/datasets/Raisin+Dataset
Rice
7
3810
1630 / 2180
2
https://archive.ics.uci.edu/ml/datasets/Rice+%28Cammeo+and+Osmancik%29
Echocardiogram
8
132
107 / 25
2
https://archive.ics.uci.edu/ml/datasets/echocardiogram
Wisconsin Diagnostic Breast Cancer
10
569
212 / 357
2
https://archive.ics.uci.edu/ml/datasets/breast+cancer+wisconsin+(diagnostic)
Loan House
11
614
422 / 192
2
@@
https://www.kaggle.com/code/sazid28/home-loan-prediction/data
Heart Failure
12
299
96 /203
2
https://archive.ics.uci.edu/ml/datasets/Heart+failure+clinical+records
Heart Disease
13
303
164 / 139
2
https://archive.ics.uci.edu/ml/datasets/heart+disease
Adult
14
32561
7841 / 24720
2
https://archive.ics.uci.edu/ml/datasets/adult
Bank Marketing
14
45211
5289 / 39922
2
https://archive.ics.uci.edu/ml/datasets/bank+marketing
Congressional Voting
16
435
267 / 168
2
https://archive.ics.uci.edu/ml/datasets/congressional+voting+records
Absenteeism
18
740
279 / 461
2
https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work
Hepatitis
19
155
32 / 123
2
https://archive.ics.uci.edu/ml/datasets/hepatitis
German
20
1000
300 / 700
2
https://archive.ics.uci.edu/ml/datasets/statlog+(german+credit+data)
Mushrooms
22
8124
4208 /3916
2
https://archive.ics.uci.edu/ml/datasets/mushroom
Credit Card
23
30000
23364 / 6636
2
https://archive.ics.uci.edu/ml/datasets/default+of+credit+card+clients
Horse Colic
26
368
232 / 136
2
https://archive.ics.uci.edu/ml/datasets/Horse+Colic
Thyroid
29
9172
2401 / 6771
2
https://archive.ics.uci.edu/ml/datasets/thyroid+disease
Cervical Cancer
31
858
55 / 803
2
https://archive.ics.uci.edu/ml/datasets/Cervical+cancer+%28Risk+Factors%29
Spambase
57
4601
1813 / 2788
2
https://archive.ics.uci.edu/ml/datasets/spambase
Iris
4
150
50 / 50 / 50
3
https://archive.ics.uci.edu/ml/datasets/iris
Balance Scale
4
625
49 / 288 / 288
3
https://archive.ics.uci.edu/ml/datasets/balance+scale
Car
6
1728
384 / 69 / 1210 / 65
4
https://archive.ics.uci.edu/ml/datasets/car+evaluation
Glass
9
214
70 / 76 / 17 / 13 / 9 / 29
6
https://archive.ics.uci.edu/ml/datasets/glass+identification
Contraceptive
9
1473
629 / 333 / 511
3
https://archive.ics.uci.edu/ml/datasets/Contraceptive+Method+Choice
Solar Flare
10
1389
1171 / 141 / 40 / 20 / 9 / 4 / 3 / 1
8
http://archive.ics.uci.edu/ml/datasets/solar+flare
Wine
12
178
59 / 71 / 48
3
https://archive.ics.uci.edu/ml/datasets/wine
Zoo
16
101
41 / 20 / 5 / 13 / 4 / 8 /10
7
https://archive.ics.uci.edu/ml/datasets/zoo
Lymphography
18
148
2 / 81 /61 /4
4
https://archive.ics.uci.edu/ml/datasets/Lymphography
Segment
19
2310
330 / 330 / 330 / 330 / 330 / 330 / 330
7
https://archive.ics.uci.edu/ml/datasets/image+segmentation
Dermatology
34
366
112 / 61 / 72 / 49 / 52 / 20
6
https://archive.ics.uci.edu/ml/datasets/dermatology
Landsat
36
6435
1533 / 703 / 1358 / 626 / 707 /1508
6
https://archive.ics.uci.edu/ml/datasets/Statlog+(Landsat+Satellite)
Annealing
38
798
8 / 88 / 608 / 60 / 34
5
https://archive.ics.uci.edu/ml/datasets/Annealing
Splice
60
3190
767 / 768 / 1655
3
https://archive.ics.uci.edu/ml/datasets/Molecular+Biology+(Splice-junction+Gene+Sequences)

### v7

@@ 2 Related Work
While optimal DTs search for a global optimum, GradTree does not necessarily pursue this. Instead, like optimization in neural networks, it aims for a local optimum that offers good generalization to test data.
We further want to emphasize that the local optima that can be reached by GradTree have a significant advantage over the local optimum of a greedy approach: While the local optimum of greedy approaches is constrained by sequentially selecting the optimal split at each node, GradTree overcomes this limitation by optimizing all parameters jointly.
@@
Axis-Aligned DT Ensembles
While existing methods for learning tree ensembles often require structural changes, such as relaxing axis-aligned splits to oblivious ones, Bruch, Pfeifer, and Guillame-Bert (2020) use input perturbation as a purely analytical procedure to approximate the gradients of the ensemble. As a result, similar to GradTree, it is possible to learn and fine-tune an axis-aligned DT ensemble. Although their method makes this feasible, the authors decided not to learn the split index but only the split threshold, in order to avoid potential changes to the model’s structure. In contrast, GradTree uses a dense representation that allows for the efficient learning of both the split threshold and the feature index to split on. This is particularly important when aiming for an interpretable DT model, as opposed to a large ensemble, since the number of possible splits is limited.
@@ 3.2 Dense Decision Tree Representation
To this end, we expand the vector $\boldsymbol{\iota}\in\mathbb{R}^{2^{d}-1}$ to a matrix $I\in\mathbb{R}^{2^{d}-1}\times\mathbb{R}^{n}$. This is achieved by one-hot encoding the feature index as $\boldsymbol{\iota}\in\mathbb{R}^{n}$ for each internal node. This adjustment is necessary for the optimization process to account for the fact that feature indices are categorical instead of ordinal. Although our matrix representation for feature selection has parallels with that proposed by Popov, Morozov, and Babenko (2019), we introduce a novel aspect: A matrix representation for split thresholds. We denote this representation as $T\in\mathbb{R}^{2^{d}-1}\times\mathbb{R}^{n}$. Instead of representing a single value for all features, we store individual values for each feature, denoted as $\boldsymbol{\tau}\in\mathbb{R}^{n}$. This modification is tailored to support the optimization process, particularly in recognizing that split thresholds are feature-specific and non-interchangeable. In essence, a viable split threshold for one feature may not be suitable for another. This adjustment acts as a memory mechanism, ensuring that a given split threshold is exclusively associated with the corresponding feature. Consequently, this refinement enhances the exploration of feature selection at every split during the training. Besides the previously mentioned advantages, using a dense DT representation allows the use of matrix multiplications for an efficient computation. Accordingly, we can reformulate the Heaviside split function (Equation 2) as
@@ 4.2 Results
Binary Multi Binary Multi Binary Multi
GradTree 054 086 0.051 0.174 0.670 0.470
DNDT 887 907 0.039 0.239 0.306 0.295
GeneticTree 007 020 0.204 0.258 0.427 0.315
DL8.5 028 029 0.202 0.260 0.371 0.331
CART 067 076 0.183 0.247 0.571 0.929
@@
ST Entmax (ours) ST Gumbel Temp. Annealing
@@ Appendix D Datasets
Blood Transfusion 4 748 178 / 570 2 https://archive.ics.uci.edu/ml/datasets/Blood+Transfusion+Service+Center
Banknote Authentication 4 1372 610 / 762 2 https://archive.ics.uci.edu/ml/datasets/banknote+authentication
Titanic 7 891 342 / 549 2 https://www.kaggle.com/c/titanic
Raisin 7 900 450 /450 2 https://archive.ics.uci.edu/ml/datasets/Raisin+Dataset
Rice 7 3810 1630 / 2180 2 https://archive.ics.uci.edu/ml/datasets/Rice+%28Cammeo+and+Osmancik%29
Echocardiogram 8 132 107 / 25 2 https://archive.ics.uci.edu/ml/datasets/echocardiogram
Wisconsin Diagnostic Breast Cancer 10 569 212 / 357 2 https://archive.ics.uci.edu/ml/datasets/breast+cancer+wisconsin+(diagnostic)
@@
Loan House 11 614 422 / 192 2 https://www.kaggle.com/code/sazid28/home-loan-prediction/data
Heart Failure 12 299 96 /203 2 https://archive.ics.uci.edu/ml/datasets/Heart+failure+clinical+records
Heart Disease 13 303 164 / 139 2 https://archive.ics.uci.edu/ml/datasets/heart+disease
Adult 14 32561 7841 / 24720 2 https://archive.ics.uci.edu/ml/datasets/adult
Bank Marketing 14 45211 5289 / 39922 2 https://archive.ics.uci.edu/ml/datasets/bank+marketing
Congressional Voting 16 435 267 / 168 2 https://archive.ics.uci.edu/ml/datasets/congressional+voting+records
Absenteeism 18 740 279 / 461 2 https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work
Hepatitis 19 155 32 / 123 2 https://archive.ics.uci.edu/ml/datasets/hepatitis
German 20 1000 300 / 700 2 https://archive.ics.uci.edu/ml/datasets/statlog+(german+credit+data)
Mushrooms 22 8124 4208 /3916 2 https://archive.ics.uci.edu/ml/datasets/mushroom
Credit Card 23 30000 23364 / 6636 2 https://archive.ics.uci.edu/ml/datasets/default+of+credit+card+clients
Horse Colic 26 368 232 / 136 2 https://archive.ics.uci.edu/ml/datasets/Horse+Colic
Thyroid 29 9172 2401 / 6771 2 https://archive.ics.uci.edu/ml/datasets/thyroid+disease
Cervical Cancer 31 858 55 / 803 2 https://archive.ics.uci.edu/ml/datasets/Cervical+cancer+%28Risk+Factors%29
Spambase 57 4601 1813 / 2788 2 https://archive.ics.uci.edu/ml/datasets/spambase
Iris 4 150 50 / 50 / 50 3 https://archive.ics.uci.edu/ml/datasets/iris
Balance Scale 4 625 49 / 288 / 288 3 https://archive.ics.uci.edu/ml/datasets/balance+scale
Car 6 1728 384 / 69 / 1210 / 65 4 https://archive.ics.uci.edu/ml/datasets/car+evaluation
Glass 9 214 70 / 76 / 17 / 13 / 9 / 29 6 https://archive.ics.uci.edu/ml/datasets/glass+identification
Contraceptive 9 1473 629 / 333 / 511 3 https://archive.ics.uci.edu/ml/datasets/Contraceptive+Method+Choice
Solar Flare 10 1389 1171 / 141 / 40 / 20 / 9 / 4 / 3 / 1 8 http://archive.ics.uci.edu/ml/datasets/solar+flare
Wine 12 178 59 / 71 / 48 3 https://archive.ics.uci.edu/ml/datasets/wine
Zoo 16 101 41 / 20 / 5 / 13 / 4 / 8 /10 7 https://archive.ics.uci.edu/ml/datasets/zoo
Lymphography 18 148 2 / 81 /61 /4 4 https://archive.ics.uci.edu/ml/datasets/Lymphography
Segment 19 2310 330 / 330 / 330 / 330 / 330 / 330 / 330 7 https://archive.ics.uci.edu/ml/datasets/image+segmentation
Dermatology 34 366 112 / 61 / 72 / 49 / 52 / 20 6 https://archive.ics.uci.edu/ml/datasets/dermatology
Landsat 36 6435 1533 / 703 / 1358 / 626 / 707 /1508 6 https://archive.ics.uci.edu/ml/datasets/Statlog+(Landsat+Satellite)
Annealing 38 798 8 / 88 / 608 / 60 / 34 5 https://archive.ics.uci.edu/ml/datasets/Annealing
Splice 60 3190 767 / 768 / 1655 3 https://archive.ics.uci.edu/ml/datasets/Molecular+Biology+(Splice-junction+Gene+Sequences)